OpenStudy (anonymous):

The scale of a town map is 1 inch to 4 miles. On the map, the park and the airport are 6 inches apart. What is the actual distance between the park and the airport?

OpenStudy (texaschic101):

1/4 = 6/x --- (1 in to 4 mi = 6 in to x mi) cross multiply (1)(x) = (4)(6) x = 24 miles

OpenStudy (anonymous):

1 inch = 4 miles 6 inches = x miles 6 * 4 = x x = 24 miles

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

you used proportions and I just used given data and we both got the same answer. There are often quite a few ways to get an answer to an equation using algebra.